IN AN ALPHA DECAY THE KE OF A PARTICLE IS 48Mev AND Q-VALUE OF REACTION IS 50Mev.THE MASS NUMBER OF MOTHER NUCLEUS IS(ASSUME THAT DAUGHTER NUCLEUS IN GROUND STATE):A)96B)100C)104D)NONE OF THESEPLS GIVE DETAIL ANALYSIS,

Grade:11

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Alpha particles are positively charged nuclei of helium, with two protons and two neutrons bound together. Alpha particles are denoted by the Greek symbol α or the chemical formula of Helium nuclei which is 42He.

The correct option is c)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

akhil samir killawala
13 years ago

I AM AWARE OF WHAT AN ALPHA PARTICLE IS!!

I WANT TO KNOW HOW TO SOLVE THE QUESTION AS YOU HAVE NOT MENTIONED ANYTHING THAT IIS GIVEN IN THE QUESTION..

PLS GIVE DETAILED ANALYSIS IF POSSIBLE

Jerry
6 years ago
For alpha decay,KE= (A-4)Q/ASubstituting,48A=(A-4)50200=2AA=1OOHope this helps%%%%%%%%%%%%%%%%%%%%%%%%%%%